# Opinion Dynamics with Multiple Adversaries

### v1

@@ 4.2. Learning the Strategic Actors with Robust Regression
Figure 7. Reconstruction error and balanced accuracy for the robust regression problem presented in Algorithm 2. The $x$-axis shows the percentage of strategic agents. The left subfigure shows the recovery error, measured as $\frac{1}{n}\sum_{i\in[n]}\left|\frac{\hat{\bm{s}}_{i}-\bm{s}_{i}}{\bm{s}_{i}}\right|$, and the right subfigure measures the balanced accuracy between the recovered $\hat{S}$ and the true $S$. To construct the confidence intervals, for each size $|S|$ of the set $S$, we draw $S$ five times randomly from the vertex set $[n]$. For the Twitter and Reddit datasets, we have used 128-dimensional Node2vec embeddings. For the Polblogs dataset we have used the community membership (which corresponds to the political orientation) of each node, such that $\bm{x}_{i}=(1,0)$ corresponds to liberal and $\bm{x}_{i}=(0,1)$ corresponds to conservative, and (the true) $\bm{s}$ is such that $\bm{s}=X\bm{v}$ for $\bm{v}=(1,-1)^{T}$. We have also provided results using 128-dimensional spectral embeddings. We have set the recovery threshold for Torrent to be $|S|/n$, and the step size to be $\eta=1/\|\bar{X}\|_{2}^{2}$ where $\bar{X}$ is the min-max normalized embedding matrix.
5. Discussion and Conclusion
In this paper, we examine how opinions evolve in social networks, where individuals adjust their publicly stated views based on interactions with others and their inherent beliefs. In our model, strategically motivated users can distort these dynamics by misrepresenting their intrinsic opinions, often to advance conflicting objectives or promote rival narratives. We analyze the Nash Equilibrium of the resulting strategic interactions and empirically show — using diverse datasets from Twitter, Reddit, and Political Blogs — that such deceptive behavior intensifies polarization, fuels disagreement, and increases equilibrium costs. Additionally, we establish worst-case guarantees on the Price of Misreporting, akin to the Price of Anarchy, and introduce scalable learning algorithms to help platforms (i) detect opinion manipulation and (ii) identify the users responsible. Our algorithms perform effectively on real-world data, suggesting how platforms might mitigate the effects of strategic opinion shaping.
We conclude with some discussion of the implications of our work, and directions for future work.
Structural Platform Interventions.
We give algorithms for platforms to detect in strategic manipulation has occurred, and who is responsible (Section 4). Having done so, a platform might seek to mitigate the effects of strategic behavior. There are multiple plausible avenues to do so.
First, they may seek to reduce degree disparities in the network by algorithmically encouraging balanced connections, such as by suggesting users connect with those who have fewer connections, or reducing the visibility of central nodes (hubs). As we saw, both the upper bound on the $\mathsf{PoM}$ and the ratio $\|\bm{s}^{\prime}\|_{2}/\|\bm{s}\|_{2}$ depend on the largest eigenvalue $\lambda_{n}$ of the Laplacian, which scales with the maximum degree. This motivates interventions to balance the degree distribution.
Second, platforms can design strategy-proof mechanisms to incentivize the agents to report their true opinions. For resource allocation games on networks, it is known that the classical Vickrey–Clarke–Groves (VCG) mechanism is susceptible to adversarial behaviors such as collusion, motivating the need for different mechanisms (Chorppath et al., 2015). In the case of social networks, platforms have unique tools such as fines or banning of accounts to modify agents’ utility functions.
Future Work
First, it is not clear which measure of centrality should be used to identify the users who are most capable of manipulating others. As can be seen from Theorem 2.2, the influence of agent $i\in S$ should depend on the other members of $S$, as well as the spectral properties of the localized Laplacian matrices $L_{j}$ for $j\in S$ and susceptibility parameters $\alpha_{k}$ for $k\in[n]$.
Second, providing $\mathsf{PoM}$ bounds where $S$ can be any set of agents constitutes another interesting research direction, especially if platforms can assume that $S$ is a small fraction of all users. As we noted after Theorem 3.1, we believe that this would require restricted invertibility analysis of the matrices determining Nash equilibria.
Third, future work might consider different models of strategic manipulation. For example, one could consider a “feedback equilibrium” model (in the sense of dynamic games (Li et al., 2024)), in which agents $i\in S$ can report arbitrary $\bm{z}_{i}^{\prime}(t)$ at each timestep $t$, rather than following Eq. (1). This flexbility may give strategic agents more power to influence outcomes.
@@ Appendix A Proofs
A.1. Proof of Proposition 4.4
@@
Appendix B Proof of Proposition 4.5
@@
Appendix C Additional Figures

### v2

@@ 4. Learning from Network Outcomes
Figure 7. Reconstruction error and balanced accuracy for the robust regression problem in Algorithm˜2 as a function of the fraction of strategic agents. The left panel reports the recovery error $\frac{1}{n}\sum_{i\in[n]}\left|\frac{\hat{\bm{s}}_{i}-\bm{s}_{i}}{\bm{s}_{i}}\right|$, while the right panel shows the balanced accuracy between the recovered set $\hat{S}$ and the true set $S$. Confidence intervals are obtained by sampling $S$ uniformly at random from $[n]$ five times for each $|S|$. For the Twitter and Reddit datasets, we use 128-dimensional Node2vec embeddings. For Polblogs, node features correspond to community membership (liberal $\bm{x}_{i}=(1,0)$, conservative $\bm{x}_{i}=(0,1)$), with intrinsic opinions given by $\bm{s}=X\bm{v}$ for $\bm{v}=(1,-1)^{\top}$. We also report results using 128-dimensional spectral embeddings. For Torrent, the recovery threshold is set to $|S|/n$ and the step size to $\eta=1/\|\bar{X}\|_{2}^{2}$, where $\bar{X}$ denotes the min–max normalized embedding matrix.
5. Managerial Insights and Discussion
5.1. Platform Governance
Our characterization of the Nash Equilibrium in multi-adversary opinion dynamics offers three strategic pillars for social media platform governance:
PoM as a Vulnerability Metric. The Price of Misreporting (PoM) quantifies the social cohesion lost to strategic dishonesty. Platform stakeholders should prioritize monitoring networks with high connectivity but low individual susceptibility, as these “viral” environments are most vulnerable to disproportionate increases in polarization from even small-scale coordinated misreporting.
Beyond Content Takedowns. Because strategic actors optimize reported opinions based on network structure, they can exert significant influence without triggering traditional platform health interventions (for example, flagging a post as fake news). Platforms must shift from individual post moderation to monitoring structural shifts and community sensitivity using “what-if” simulations of the Nash Equilibrium.
Structural Friction. To narrow the gap between reported and intrinsic opinions, platforms can effectively increase the sincerity cost by implementing friction mechanisms. Increasing transparency of historical stances and cross-network consistency makes maintaining strategic “lies” more costly, disincentivizing extreme misreporting.
5.2. Conclusion
We introduced a game-theoretic framework that replaces the "single bad actor" assumption with a realistic model of multiple, competing adversaries. We proved the existence of a unique Nash Equilibrium and introduced the Price of Misreporting to bound social inefficiency. Our empirical results across Reddit, Twitter, and Political Blogs datasets demonstrate that multi-agent competition often heightens global disagreement rather than neutralizing it. These findings provide a mathematical foundation for a new generation of intervention tools that prioritize the structural integrity of social graphs over keyword-based moderation. Future research should investigate how these strategic behaviors evolve alongside adaptive network topologies.
Acknowledgements.
The authors would like to thank Yanbang Wang, Yixuan He, Jon Kleinberg, Giannis Fikioris, and Olivia Sheng for their insightful comments on the current manuscript.
@@ Appendix A Proofs
Proof of Theorem˜2.2
$\nabla_{\bm{s}^{\prime}}c_{i}(\bm{s}^{\prime})$ $=(1-\alpha_{i})\sum_{j\sim i}w_{ij}2(B^{T}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}B)(\bm{s}^{\prime})$
$+\alpha_{i}(2B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B\bm{s}^{\prime}-2B^{T}\bm{e}_{i}\bm{e}_{i}^{T}\bm{s}),$
$\nabla_{\bm{s}^{\prime}}^{2}c_{i}(\bm{s}^{\prime})$ $=2(1-\alpha_{i})B^{T}\bigg[\sum_{j\sim i}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}\bigg]B+2\alpha_{i}B^{T}\bm{e}_{i}\bm{e}_{i}^{T}B.$
Let $L_{i}\in\mathbb{R}^{n\times n}$ be:
$L_{i}:=\sum_{j\sim i}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}.$
$0$ $=\frac{1}{2}\frac{\partial}{\partial\bm{s}_{i}^{\prime}}c_{i}(\bm{s}^{\prime})$
$=\bm{e}_{i}^{T}(1-\alpha_{i})B^{T}\bigg[\sum_{j\sim i}w_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{T}\bigg]B\bm{s}^{\prime}$
$+\bm{e}_{i}^{T}\alpha_{i}(B^{T}\bm{e}_{i}\bm{e}_{i}^{T}Bs^{\prime}-B^{T}\bm{e}_{i}\bm{e}_{i}^{T}\bm{s})$
The above display gives the solution for $\bm{s}_{i}^{\prime}$ in terms of all entries of $s^{\prime}$. Assembling the critical points into a linear system, we obtain precisely that for all $i\in S$, $\bm{e}_{i}^{T}\mathbb{T}_{i}\bm{s}^{\prime}=\bm{y}_{i}$. Since $\bm{s}_{j}^{\prime}=\bm{s}_{j}$ for $j\not\in S$, the overall linear system describes the Nash equilibria.
Proof of Proposition˜4.4
@@
Proof of Theorem 3.1
(11) $Q=BLB+\tilde{\alpha}(I-2B+B^{2})$
[…]
Next, let $\textrm{diag}(B)$ be the diagonal matrix with entries $B_{ii}$ and $\widetilde{\textrm{diag}(B)}$ be the diagonal matrix with elements $B_{ii}$. It is easy to show that $\bm{s}^{\prime}=(1/\tilde{\alpha})B^{-1}\widetilde{\textrm{diag}(B)}\bm{s}$ and $\bm{z}^{\prime}=(1/\tilde{\alpha})\widetilde{\textrm{diag}(B)}\bm{s}$, which similarly implies (after algebraic operations) that $C(\bm{z}^{\prime})/(1-\alpha)=\bm{s}^{T}Q^{\prime}\bm{s}$ where:
$Q^{\prime}$ $:=\frac{1}{\tilde{\alpha}^{2}}\widetilde{\textrm{diag}(B)}L\widetilde{\textrm{diag}(B)}+\frac{1}{\tilde{\alpha}}B^{-1}\left(\widetilde{\textrm{diag}(B)}\right)^{2}B^{-1}$
$-2\frac{1}{\tilde{\alpha}}B^{-1}\left(\widetilde{\textrm{diag}(B)}\right)^{2}+\frac{1}{\tilde{\alpha}}\left(\widetilde{\textrm{diag}(B)}\right)^{2}$
(12) $\|\widetilde{\textrm{diag}(B)}\|_{2}$ $=\max_{i}B_{ii}\leq\|B\|_{2}=1,$
By the triangle inequality, the Cauchy-Schwarz inequality, and Equations˜12 and 13, we have that:
$\|Q^{\prime}\|_{2}$ $\leq\frac{1}{\tilde{\alpha}^{2}}\|L\|_{2}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}+\frac{1}{\tilde{\alpha}}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}\|B^{-1}\|_{2}^{2}$
$+\frac{2}{\tilde{\alpha}}\|B^{-1}\|_{2}+\frac{1}{\tilde{\alpha}}\left(\|\widetilde{\textrm{diag}(B)}\|_{2}\right)^{2}$
Finally, we can simplify:
Extension to heterogeneous $\alpha_{i}$
Corollary A.1 (Price of Misreporting for Heterogeneous Susceptibility).
Note that $C(\bm{z}^{\prime})\leq(1-\alpha_{\min})(\bm{z}^{\prime})^{T}L\bm{z}^{\prime}+\alpha_{\max}\|\bm{z}^{\prime}-\bm{s}\|_{2}^{2}=\overline{C}(\bm{z}^{\prime})$, and $C(\bm{z})\geq(1-\alpha_{\max})\bm{z}^{T}L\bm{z}+\alpha_{\min}\|\bm{z}-\bm{s}\|_{2}^{2}=\underline{C}(\bm{z})$ where $\alpha_{\min}=\min_{i\in[n]}\alpha_{i}$, and $\alpha_{\max}=\max_{i\in[n]}\alpha_{i}$. Then, the same analysis of Theorem˜3.1 can be applied, since $\overline{C}(\bm{z}^{\prime})/\underline{C}(\bm{z})\geq C(\bm{z}^{\prime})/C(\bm{z})$.
Proof of Proposition˜4.2
Let $\widehat{\bm{s}^{\prime}}$ be as in Algorithm 2, and $\bm{y}=\widehat{\bm{s}^{\prime}}$. Notice $\bm{y}=\bm{s}+(\bm{s}^{\prime}-\bm{s})+(\widehat{\bm{s}^{\prime}}-\bm{s}^{\prime})$. Let $\bm{w}:=(\bm{s}^{\prime}-\bm{s})$ be the corruption vector due to strategic negotiations and $\bm{r}=(\widehat{\bm{s}^{\prime}}-\bm{s}^{\prime})$ be the residual vector due to least-squares regression. We claim that $\bm{r}=\bm{0}$, because $A^{-1}$ is full rank and ((IA)L+A)((I-A)L+A) is full rank, so $\widehat{\bm{s}^{\prime}}=A^{-1}((I-A)L+A)\bm{z}^{\prime}=\bm{s}^{\prime}$.
Therefore, letting $T=C^{\prime}(\log n)^{2}$ for large enough constant $C^{\prime}>0$, we see that $\|\hat{\bm{v}}-\bm{v}\|_{2}\leq O(n^{-\omega(1)})$. Hence $\|\hat{\bm{s}}-\bm{s}\|_{2}=\|X\hat{\bm{v}}-X\bm{v}\|_{2}\leq\|X\|_{2}\cdot n^{-\omega(1)}$. Now, let $\bm{u}=\hat{\bm{s}}-\bm{s}^{\prime}$. If $i\in[n]\setminus S$, then $\lvert\bm{u}_{i}\rvert\leq\|X\|_{2}\cdot n^{-\omega(1)}$. On the other hand for $j\in S$, $\lvert\bm{u}_{j}\rvert\geq\lvert\bm{s}_{j}-\bm{s}_{j}^{\prime}\rvert-\|X\|_{2}n^{-\omega(1)}$. Therefore the top-$|S|$ entries of $\bm{u}$ recover $S$.
Proof of Proposition˜4.5
@@
Appendix B Additional Figures